# Ov3R: Open-Vocabulary Semantic 3D Reconstruction from RGB Videos

### v1

@@ 1 CLIP3R
Subsequently, the concatenated features $F_{cat}$ ($F_{cat}^{f},F_{cat}^{s}\in F_{cat}$) and CLIP features are processed through cross-attention (CA) to obtain the features $F_{atten-cd}$, guided by the spatial structure in DINO features. We sum up CLIP features, DINO features, and the refined CLIP features to represent the descriptors of $I_{full}$ and $I_{seg}$ as:
$d_{full}=F_{clip}^{f}+F_{cat}^{f}+softmax(\frac{F_{cat}^{f}F_{clip}^{f}}{F_{clip}^{f}})F_{clip}^{f}$ (8)
$d_{seg}=F_{clip}^{s}+F_{cat}^{s}+softmax(\frac{F_{cat}^{s}F_{clip}^{s}}{F_{clip}^{s}})F_{clip}^{s}$ (9)
where $d_{full}$ and $d_{seg}$ denote the descriptors of the full frame $I_{full}$ and the image crop containing the 2D segment (w/ background) $I_{seg}$, respectively. In the third branch, we introduce a CLIP-like distilled 3D point encoder (Hegde, Valanarasu, and Patel 2023) to extract geometric features. This 3D encoder is pre-trained on triplets of pointclouds, corresponding images, and text using natural language supervision. The 3D features obtained by this distilled encoder naturally align with the CLIP latent space. Accordingly, we first obtain the 3D segment $P_{oseg}$ corresponding to the 2D segment (w/o background) $I_{oseg}$. Then, the 3D features $F_{point}^{os}$, processed by a linear layer, are fed into cross-attention with CLIP features $F_{clip}^{os}$ to learn semantic-geometric relationships. We add the geometry-informed features to the CLIP features to represent the descriptor of $I_{oseg}$:
$d_{oseg}=F_{clip}^{os}+softmax(\frac{F_{point}^{os}F_{clip}^{os}}{F_{clip}^{os}})F_{clip}^{os}$ (10)
where $d_{oseg}$ denotes the descriptor of the image crop only with the 2D segment $I_{oseg}$. We employ the same weighted merging strategy from (Martins, Oswald, and Civera 2024) to combine these three features from different levels:
$d=\sum^{i}w_{i}\odot d_{i},i=full,seg,oseg$ (11)
where $d$ is our 2D-3D fused descriptor. $\odot$ denotes the Hadamard product. The weights $w_{i}$ are obtained by a shallow model comprising a cross-attention layer, an MLP, and a softmax layer. We pre-trained our 2D-3D fusion model by minimizing the sigmoid cosine similarity loss:
$L_{sim}=-\frac{1}{\left|B\right|}\sum_{i}^{\left|B\right|}\sum_{j}^{\left|B\right|}log\left(\frac{1}{1+exp(z_{ij}(-kd_{i}\cdot t_{j}+b))}\right)$ (12)
@@ 4 Experiments
Datasets. For the 3D reconstruction task, we train CLIP3R on ScanNet++ (Yeshwanth et al. 2023), Aria Synthetic Environments (Avetisyan et al. 2024), and CO3D-v2 (Reizenstein et al. 2021), which provide diverse scenarios and objects from both real-world and synthetic scenes. We evaluate 3D reconstruction performance on Replica (Straub et al. 2019) and 7Scenes (Shotton et al. 2013). For the open-vocabulary 3D segmentation task, we train the 2D-3D OVS module on ScanNet++ and evaluate it on Replica and ScanNetv2 (Dai et al. 2017). Additionally, we report camera tracking results from CLIP3R on both Replica and 7Scenes.
@@
1 3D Reconstruction
7Scenes. Table 2 reports the evaluation on the 7Scenes dataset, carried out by uniformly sampling one-twentieth of the frames from each test sequence following (Liu et al. 2025). As shown in the table, on average, Ov3R surpasses all of the baselines in both accuracy and completion, confirming the trends already highlighted on the Replica dataset.
Figure 6: Visualization of open-vocabulary 3D queries on the Replica. We visualize two queries from ”Head” and ”Common” in room0. The similarities of queries are shown from low (blue) to mid (green), and high (yellow). OVO yields wrong ”Pillows” and the missing ”Vase”, while Open3DIS also produces wrong ”Pillows” and low-similarity ”Vase”.
2 Camera Tracking
Although 3R models do not explicitly estimate trajectories, camera poses can be derived from their predicted scene points. Following (Wang et al. 2024b), we extract poses by running PnP-RANSAC (Fischler and Bolles 1981) from the predicted 3D points, given the known camera intrinsics of each frame. Accordingly, we assess the tracking performance of Ov3R on the Replica and 7Scenes datasets. As shown in Table 1 and Table 2, Ov3R outperforms the other two real-time 3R-based frameworks, Spann3R and SLAM3R, narrowing the gap between this novel family of methods and the much slower dense SLAM systems.
3 Open-Vocabulary 3D Semantic Segmentation
@@
Replica. We run experiments on the 8 standard scenes of Replica, annotated with 51 different semantic classes, and evaluate the results directly in 3D space rather than by projecting semantic masks onto 2D images (Kerr et al. 2023). Following (Engelmann et al. 2024), these labels are divided into three categories (”Head”, ”Common”, ”Tail”) according to their frequency in the dataset. Thanks to its modular design, we compare Ov3R with both offline approaches and the recent online method, OVO (Martins, Oswald, and Civera 2024). In Table 3, Ov3R outperforms all baselines on mIoU and mAcc across All labels on average, achieving the best trade-off across all categories. Although Open3DIS achieves better performance on ”Head” labels, it fails on less frequent objects. As shown in Figure 6, Ov3R accurately identifies the queried semantic classes, whereas OVO and Open3DIS produce inaccurate segmentation or miss objects entirely.

### v2

@@ 3.1 CLIP3R
These features are cross-attended to original CLIP3R features at both levels independently (CA block in the figure). The resulting features are finally summed with the previous concatenated features and CLIP3R features ($\oplus$ in the figure).
$\mathbf{D}_{\text{scene}}=\mathbf{F}_{\text{CLIP3R}}^{\text{scene}}+\mathbf{F}_{cat}^{\text{scene}}+\texttt{softmax}(\frac{\mathbf{F}_{\text{CLIP3R}}^{\text{scene}}\cdot\mathbf{F}_{cat}^{\text{scene }T}}{\sqrt{d}})\cdot\mathbf{F}_{cat}^{\text{scene}}$ (9)
$\mathbf{D}_{\text{inst}}=\mathbf{F}_{\text{CLIP3R}}^{\text{inst}}+\mathbf{F}_{cat}^{\text{inst}}+\texttt{softmax}(\frac{\mathbf{F}_{\text{CLIP3R}}^{\text{inst}}\cdot\mathbf{F}_{cat}^{\text{inst }T}}{\sqrt{d}})\cdot\mathbf{F}_{cat}^{\text{inst}}$ (10)
In parallel, we introduce a CLIP-like distilled 3D point encoder [hegde2023clip] to extract geometric features. This 3D encoder is pre-trained on triplets of point clouds, corresponding images, and text using natural language supervision. The 3D features obtained by this distilled encoder naturally align with the CLIP latent space. These 3D features are then processed by a linear layer and cross-attended with masked object-level CLIP3R features to learn semantic-geometric relationships, and finally summed to these latter again similarly to what done in the DINO branch.
$\mathbf{D}_{\text{mask}}=\mathbf{F}_{\text{CLIP3R}}^{\text{mask}}+\texttt{softmax}(\frac{\textbf{F}_{\text{CLIP3R}}^{\text{mask}}\cdot\mathbf{F}_{\text{3D-CLIP}}^{T}}{\sqrt{d}})\cdot\mathbf{F}_{\text{3D-CLIP}}$ (11)
Finally, we employ the weighted merging strategy from [martins2024ovo] to combine these three features from different levels:
$\mathbf{D}=w_{\text{scene}}\odot\mathbf{D}_{\text{scene}}+w_{\text{inst}}\odot\mathbf{D}_{\text{inst}}+w_{\text{mask}}\odot\mathbf{D}_{\text{mask}}$ (12)
where $\mathbf{D}$ is our 2D-3D fused descriptor. $\odot$ denotes the Hadamard product. The weights $w_{i}$ are obtained by a shallow model comprising a cross-attention layer, an MLP, and a softmax layer. We pre-trained our 2D-3D fusion model by minimizing the sigmoid cosine similarity loss:
$\mathcal{L}_{\text{sim}}=-\frac{1}{\left|B\right|}\sum_{i}^{\left|B\right|}\sum_{j}^{\left|B\right|}log\left(\frac{1}{1+e^{z_{ij}(-kd_{i}\cdot t_{j}+b)}}\right)$ (13)
@@ 4 Experiments
Datasets. For the 3D reconstruction task, we follow [liu2025slam3r] and train CLIP3R on ScanNet++ [yeshwanth2023scannet++], Aria Synthetic Environments [avetisyan2024scenescript], and CO3D-v2 [reizenstein2021common], which provide diverse scenarios and objects from both real-world and synthetic scenes. We evaluate 3D reconstruction performance on Replica [straub2019replica] and 7Scenes [shotton2013scene]. For the open-vocabulary 3D segmentation task, we train the 2D-3D OVS module on ScanNet++ and evaluate it on Replica and ScanNetv2 [dai2017scannet], following [martins2024ovo]. Additionally, we report camera tracking results from CLIP3R on both Replica and 7Scenes.
@@
4.1 3D Reconstruction
7Scenes. Table 2 reports the evaluation on the 7Scenes dataset, carried out by uniformly sampling one-twentieth of the frames from each test sequence following [liu2025slam3r]. As shown in the table, on average, Ov3R surpasses all of the baselines in both accuracy and completion, confirming the trends already highlighted on the Replica dataset.
@@
Figure 6: Visualization of open-vocabulary 3D queries on the Replica. We visualize two queries from ”Head” and ”Common” in room0. The similarities of queries are shown from low (blue) to mid (green), and high (yellow). OVO-mapping yields wrong ”Pillows” and the missing ”Vase”, while Open3DIS also produces wrong ”Pillows” and low-similarity ”Vase”.
4.2 Camera Tracking
Although 3R models do not explicitly estimate trajectories, camera poses can be derived from their predicted scene points. Following [wang2024dust3r], we extract poses by running PnP-RANSAC [fischler1981random] from the predicted 3D points, given the known camera intrinsics of each frame. Accordingly, we assess the tracking performance of Ov3R on the Replica and 7Scenes datasets. As shown in Table 1 and Table 2, Ov3R outperforms the other two real-time 3R-based frameworks, Spann3R and SLAM3R, narrowing the gap between this novel family of methods and the much slower dense SLAM systems.
4.3 Open-Vocabulary 3D Semantic Segmentation
@@
Replica. We run experiments on the 8 standard scenes of Replica, annotated with 51 different semantic classes, and evaluate the results directly in 3D space rather than by projecting semantic masks onto 2D images [kerr2023lerf]. Following [engelmann2024opennerf], these labels are divided into three categories (”Head”, ”Common”, ”Tail”) according to their frequency in the dataset. Thanks to its modular design, we compare Ov3R with both offline approaches and the recent online method, OVO [martins2024ovo]. In Table 3, Ov3R outperforms all baselines on mIoU and mAcc across All labels on average, achieving the best trade-off across all categories. Although Open3DIS achieves better performance on ”Head” labels, it fails on less frequent objects. As shown in Figure 6, Ov3R accurately identifies the queried semantic classes, whereas OVO and Open3DIS produce inaccurate segmentation or miss objects entirely.